During the above window we will be replacing server components that we have identified as degraded. This will affected “Dedicated VoIP” on the affected hardware for a short period of time. We apologise for any inconvenience caused.
UPDATE01 – 12:15
This work is completed and VM PBXs restored.
We have been receiving reports of poor quality affecting some FTTC connections. We are currently working to identify the issue.
If you are experiencing problems please call in to report the issue so we can continue to collect affected circuits.
UPDATE01 – 09:39
We have collated a number of lines and reported these to our wholesale provider who have suggested a possible hardware fault in London within there core switches. It has not been confirmed these are at fault, however we are connected to these devices.
UPDATE02 – 10:00
Ticket updated to reflect Packet Loss over call quality. Our NOC engineers are currently on a call with our supplier for an update. We apologise for the continued delay.
We are seeing packet loss clear and services are starting to recover. We have not yet received an official clear
Services remain stable. we have been advised the problem was with a 10Gbps bearer that provides a number of services from Openreach POPs. Our network monitoring showed this problem starting at 23:15 last night so we have logged a complaint with our wholesale provider to advise why this was not picked up and resolved over night prior to it becoming service affecting this AM.
Once again we apologise for any inconvenience caused and will provide an RFO on request once made available to ourselves.